# Global and Local Prompts Cooperation via Optimal Transport for Federated Learning

### v1

@@ Abstract
Prompt learning in pretrained visual-language models has shown remarkable flexibility across various downstream tasks. Leveraging its inherent lightweight nature, recent research attempted to integrate the powerful pretrained models into federated learning frameworks to simultaneously reduce communication costs and promote local training on insufficient data. Despite these efforts, current federated prompt learning methods lack specialized designs to systematically address severe data heterogeneities, e.g., data distribution with both label and feature shifts involved. To address this challenge, we present Federated Prompts Cooperation via Optimal Transport (FedOTP), which introduces efficient collaborative prompt learning strategies to capture diverse category traits on a per-client basis. Specifically, for each client, we learn a global prompt to extract consensus knowledge among clients, and a local prompt to capture client-specific category characteristics. Unbalanced Optimal Transport is then employed to align local visual features with these prompts, striking a balance between global consensus and local personalization. Extensive experiments on datasets with various types of heterogeneities have demonstrated that our FedOTP outperforms the state-of-the-art methods.
@@ 1 Introduction
Recently, vision-language pre-trained models like Contrastive Language-Image Pretraining (CLIP) [55] have shown potential in learning robust and versatile representations suitable for various image distributions, aligning with the objectives of federated learning. However, the substantial communication overhead between the server and clients renders training CLIP in a federated learning environment. Additionally, overfitting concerns may arise when large-scale models are trained with limited client data. Prompt learning [42, 73, 72] provides a flexible way to adapt pre-trained models to downstream tasks by training only additional parameters. This enables prompts to capture task-specific information while guiding the fixed model’s performance. Leveraging its lightweight nature, prior research [71, 26] has explored the integration of prompt learning into federated learning frameworks to overcome the problems outlined above.
@@
To resolve these limitations, we propose simultaneously learning both a shared global prompt and a personalized local prompt for each client in the local training phase. After local training, the local prompt is retained locally, while the global prompt is transmitted to the server to aggregate with prompts from other clients using FedAvg [48]. In this manner, the client owns the capacity to acquire consensus knowledge among clients from the global prompt, while also being able to discern client-specific user traits through the local prompt.
To further achieve a balance between global consensus and local personalization, we introduce Federated Prompts Cooperation via Optimal Transport (FedOTP). FedOTP utilizes Optimal Transport (OT) [31] to align local visual features with both global and local textual features through an adaptive transport plan, promoting fine-grained matching across vision and language modalities and strengthening collaboration between the global and local prompts. The adaptive OT transport plan can provide resilience to visual misalignment and effective adaptation to feature shifts. It’s worth noting that the standard OT formulation imposes two hard equality constraints on the transport plan, leading to each image patch being assigned to prompts. This may potentially cause prompts to capture some class-irrelevant information from the image and consequently influence the final results. To mitigate this, we consider employing unbalanced OT by relaxing one of the equality constraints, allowing prompts to focus solely on the most relevant image patches rather than the entire content of the image. For an efficient solution, we apply a fast implementation of Dykstra’s algorithm [15] in our FedOTP, enabling swift convergence and focusing on the core area of the image during iterations.
@@
  • We are the first to explore the mechanism of prompts’ cooperation in federated learning where severe data heterogeneity is present. More precisely, we train a global prompt to learn consensus information among clients, and a local prompt to capture client-specific category characteristics at the same time.
  • We propose FedOTP, a federated learning framework utilizing unbalanced OT to enhance the cooperation between global and local prompts. By aligning the local visual features and these two textual prompts, our FedOTP can effectively deal with severe data heterogeneity.
  • We conducted extensive experiments on widely adopted datasets in various data heterogeneity with domain discrepancy and imbalanced class distribution, and the significant result improvement verifies the superiority of our FedOTP. In addition, we demonstrated the ability of FedOTP to balance consensus and local personalization through visualizations.
@@ 2.1 Personalized Federated Learning
Personalized federated learning (PFL) is a highly regarded research field because of its potential to address statistical and systemic heterogeneity across clients. Various approaches have been proposed in prior research to achieve PFL. The most common method involves the inclusion of regularization terms in the loss function [38, 39, 61], and fine-tuning the global model on clients’ local datasets [67, 46, 16, 33]. Additionally, some works focus on explicitly seeking a trade-off between the global model and the local models [27, 41, 47, 7]. To enhance adaptability to diverse data distributions, certain techniques have delved into clustering methods for client grouping [56, 30, 65]. Leveraging the relationships and data distribution among clients, methods like FedPAC [68], and FedDisco [70] introduce novel weighted aggregation techniques to promote intensive collaboration among similar clients. Furthermore, some researchers have explored the decomposition of model parameters into base layers and personalized layers. For instance, FedPer [1], FedRep [11], and FedBABU [51] learn personalized classifier heads locally while sharing the base layers, and FedTP [35] learns personalized self-attention layers for each client.
@@ 2.3 Optimal Transport
Initially developed as a solution to optimize the cost of moving multiple items concurrently, Optimal Transport (OT) [31] has gained significant attention in the machine learning and computer vision community. To accelerate the convergence and efficiently deal with large-scale problems, [12] introduced Sinkorn’s algorithm [58] for computing an approximate transport coupling with entropic regularization. Classical OT lacks the flexibility for partial displacement in the transport plan. Unbalanced OT addresses this by relaxing the equality constraint, and by incorporating soft penalties based on Kullback-Leibler divergence [22], it can be efficiently solved through the generalized Sinkhorn’s algorithm [9]. Given its remarkable ability in distribution matching, OT has been applied in various theoretical and practical tasks, including domain adaptation [13, 18, 4], learning with noisy labels [20, 5], causal discovery [37, 63, 64], federated learning [17, 8] and so on. In prompt learning field, PLOT [6] proposes to learn multiple prompt sets for diverse contextual representations and use OT to align the features of vision and language modalities. Different from PLOT, we employ unbalanced OT to enhance the cooperation between global and local prompts by relaxing one of the equality constraints, which allows prompts to concentrate exclusively on the most relevant image patches.
@@ 5 Experiments
In this section, we conduct comprehensive experiments to numerically evaluate our FedOTP in the scenarios of heterogeneous data distribution.
@@ 5.1 Experimental Setup
Datasets and Data Heterogeneity. We evaluated the performance of our framework on nine public benchmark datasets with different types of heterogeneity, including label shift and feature shift. To investigate label shift heterogeneity, we selected two types of datasets. Following previous research [26, 25], we utilized five representative visual classification datasets to simulate datasets with limited samples: Food101 [3], DTD [10], Caltech101 [19], Flowers102 [50], and OxfordPets [52]. Referring to these datasets as the CLIP dataset for convenience, we utilized a Pathological setting by randomly allocating a distinct number of non-overlapping classes to each client. We also employed two image benchmark datasets: CIFAR-10, and CIFAR-100 [34]. We considered the Dirichlet Distribution as introduced in [57, 35] where the datasets are partitioned randomly among clients using a symmetric Dirichlet distribution with $\alpha=0.3$. For feature shift heterogeneity, we evaluated our framework on the following two datasets with multiple domains: DomainNet [54] with 6 domains, and Office-Caltech10 [24] with 4 domains. In line with prior studies [40, 62], each client participating in the federated learning system is assigned data from one of these distinct domains. A detailed introduction of each dataset and details about the Non-IID settings can be found in the Appendix Section B.1.
Baselines. We compared FedOTP with three kinds of baselines: (1) Local training methods: (i) Zero-shot CLIP [55] with hand-crafted text prompt templates; (ii) CoOp [73] with learnable prompt vectors trained on each client locally. (2) Existing prompt-based federated learning: (i) PromptFL [26] using a unified prompt learned across clients using FedAvg [48]; (ii) pFedPrompt [25] learning a shared prompt with personalized visual attention modules for each client. (3) Four adapted baseline methods derived from traditional PFL techniques, including PromptFL+FT [23], PromptFL+FedProx [38], PromptFL+FedPer [1] and PromptFL+FedAMP [30], as introduced in [25].
Implementation Details. To simulate federated learning in various scenarios, we consider the following two settings: (1) $n=10$ clients with a full $r=100\%$ partition, (2) $n=100$ clients with a $r=10\%$ partition. We employ SGD optimizer with a learning rate $lr=0.001$ and local epoch $R=5$ for CLIP datasets while $R=1$ for other cases. The communication round is set to $T=10$ for CLIP datasets with $10$ clients and $T=150$ for CIFAR-10/CIFAR-100 datasets with $100$ clients. We present the results using two representative backbones, ResNet50 [28] and ViT_B16 [14], defaulting to ViT_B16 if not explicitly specified. All experiments are conducted with Pytorch [53] on NVIDIA A40 GPUs. More implementation details can be found in the Appendix Section B.2.
@@ 5.2 Performance Evaluation
Model Evaluation on Label Shifts. We first measured the performance of FedOTP against baselines on datasets with label shifts. The experimental results on CLIP datasets and CIFAR-10/CIFAR-100 datasets are summarized in Table 1 and Table 3. For easy comparison, Table 1 reports results utilizing ResNet50 as the backbone, maintaining consistency with [25]. As shown in Table 1, our FedOTP outperforms state-of-the-art algorithms by a large margin across all datasets, which confirms the effectiveness of our Global-Local prompt cooperation mechanism to handle label shift scenarios. Remarkably, while both PromptPer (which splits the learnable prompt vector into “base+personalized” vectors) and pFedPrompt (utilizing a shared prompt with a personalized attention module in the vision modal) experience significant declines when datasets are altered, FedOTP exhibits slight fluctuations. This verifies the robustness of our method across diverse scenarios. Table 3 shows the results of our FedOTP and benchmark methods on CIFAR-10/CIFAR-100 datasets under Dirichlet setting over $100$ clients with $10\%$ partition. Even in this scenario with Dirichlet settings and a large number of clients, FedOTP consistently outperforms the baseline methods, further highlighting the superiority of our approach.
@@ 10                   Calculate the cost matrix $C_{k}=1-G_{m}^{\top}H_{k}$ of each class;
CIFAR10 [34] Image Classification 10 50,000 10,000 1
CIFAR100 [34] Image Classification 100 50,000 10,000 1
@@ B.2 Implementation Details
All input images across datasets are resized to $224\times 224$ pixels and further divided into $14\times 14$ patches with a dimension of $768$. Regarding the hyperparameters for solving OT, we set the entropic regularization weight in Problem Eq. (8) as $\lambda=0.1$ for all datasets. The maximum iteration number $n$ for Algorithm 1 is set to $100$, and we implement early stopping when the absolute update value $\Delta_{v}$ is less than $0.001$. For the setting of learnable prompts, the length of prompt vectors $s$ is set to $16$ with a dimension of $512$, “end” token position, and “random” initialization. Batch sizes are set to $32$ for training and $100$ for testing.

### v2

@@ Abstract
Prompt learning in pretrained visual-language models has shown remarkable flexibility across various downstream tasks. Leveraging its inherent lightweight nature, recent research attempted to integrate the powerful pretrained models into federated learning frameworks to simultaneously reduce communication costs and promote local training on insufficient data. Despite these efforts, current federated prompt learning methods lack specialized designs to systematically address severe data heterogeneities, e.g., data distribution with both label and feature shifts involved. To address this challenge, we present Federated Prompts Cooperation via Optimal Transport (FedOTP), which introduces efficient collaborative prompt learning strategies to capture diverse category traits on a per-client basis. Specifically, for each client, we learn a global prompt to extract consensus knowledge among clients, and a local prompt to capture client-specific category characteristics. Unbalanced Optimal Transport is then employed to align local visual features with these prompts, striking a balance between global consensus and local personalization. By relaxing one of the equality constraints, FedOTP enables prompts to focus solely on the core regions of image patches. Extensive experiments on datasets with various types of heterogeneities have demonstrated that our FedOTP outperforms the state-of-the-art methods.
@@ 1 Introduction
Recently, vision-language pre-trained models like Contrastive Language-Image Pretraining (CLIP) [60] have shown potential in learning robust and versatile representations suitable for various image distributions, aligning with the objectives of federated learning. However, the substantial communication overhead between the server and clients renders training CLIP in federated learning frameworks. Besides, overfitting concerns may arise when large-scale models are trained with limited client data. Prompt learning [46, 78] provides a flexible way to adapt pre-trained models to downstream tasks by training only additional parameters. This enables prompts to capture task-specific information while guiding the fixed model’s performance. Leveraging its lightweight nature, prior research [76, 27] has explored the integration of prompt learning into federated learning to overcome the problems outlined above.
@@
To resolve these limitations, we propose simultaneously learning both a shared global prompt and a personalized local prompt for each client in the local training phase. After local training, the local prompt remains on the client side, while the global prompt is transmitted to server to aggregate with prompts from other clients. In this manner, the client owns the capacity to acquire consensus knowledge among clients from the global prompt, while also being able to discern client-specific user traits through the local prompt.
To further achieve a balance between global consensus and local personalization, we introduce Federated Prompts Cooperation via Optimal Transport (FedOTP). FedOTP utilizes Optimal Transport (OT) [33] to align local visual features with both global and local textual features through an adaptive transport plan, promoting fine-grained matching across vision and language modalities and strengthening collaboration between the global and local prompts. The adaptive OT transport plan can provide resilience to visual misalignment and effective adaptation to feature shifts. It’s worth noting that the standard OT imposes two hard equality constraints on the transport plan, leading to each image patch being assigned to prompts. This may potentially cause prompts to capture some class-irrelevant information from the image and consequently influence the final results. To mitigate this, we consider employing unbalanced OT by relaxing one of the equality constraints, allowing prompts to focus solely on the most relevant image patches rather than the entire content of the image. For an efficient solution, we apply a fast implementation of Dykstra’s algorithm [16] in our FedOTP, enabling swift convergence and focusing on the core area of the image during iterations.
@@
  • We are the first to explore the mechanism of prompts’ cooperation in federated learning where severe data heterogeneity is present. More precisely, we train both a global prompt for consensus across clients and a local prompt to capture client-specific category traits simultaneously.
  • We propose FedOTP, a federated learning framework utilizing unbalanced OT to enhance the cooperation between global and local prompts. Through unbalanced OT, we align local visual features with textual prompts while enable prompts to focus solely on the critical image patches.
  • We conducted extensive experiments on widely adopted datasets in various data heterogeneity with feature shifs and label shifts, and significant result improvement verifies the superiority of our FedOTP. In addition, we demonstrated the ability of FedOTP to balance consensus and local personalization through visualizations.
@@ 2.1 Personalized Federated Learning
Personalized federated learning (PFL) is a highly regarded research field because of its potential to address statistical and systemic heterogeneity across clients. Various approaches have been proposed in prior research to achieve PFL. The most common method involves the inclusion of regularization terms in the loss function [42, 43, 66], and fine-tuning the global model on clients’ local datasets [72, 50, 17, 35]. Additionally, some works focus on explicitly seeking a trade-off between the global model and the local models [28, 45, 51, 8]. To enhance adaptability to diverse data distributions, certain techniques have delved into clustering methods for client grouping [61, 32, 70]. Leveraging the relationships and data distribution among clients, methods like FedPAC [73], and FedDisco [75] introduce novel weighted aggregation techniques to promote intensive collaboration among similar clients. Furthermore, some researchers have explored the decomposition of model parameters into base layers and personalized layers. For instance, FedPer [1], FedRep [12], and FedBABU [55] learn personalized classifier heads locally while sharing the base layers, and FedTP [39] learns personalized self-attention layers for each client. The importance of PFL was pointed out theoretically by [31].
@@ 2.3 Optimal Transport
Initially developed as a solution to optimize the cost of moving multiple items concurrently, Optimal Transport (OT) [33] has gained significant attention in the machine learning and computer vision community. To accelerate the convergence and efficiently deal with large-scale problems, [13] introduced Sinkorn’s algorithm [63] for computing an approximate transport coupling with entropic regularization. Unbalanced OT relaxes the equality constraint of classical OT, allowing for partial displacement in the transport plan. It can be efficiently solved using the generalized Sinkhorn’s algorithm [10] by incorporating soft penalties based on Kullback-Leibler divergence [23] or parallel algorithm [37]. Given its remarkable ability in distribution matching, OT has been applied in various theoretical and practical tasks, including domain adaptation [14, 19, 5, 38], learning with noisy labels [21, 6], causal discovery [41, 68, 69], federated learning [18, 9], outlier detection [59] and so on. In prompt learning field, PLOT [7] proposes to learn multiple prompt sets for diverse contextual representations and use OT to align the features of vision and language modalities. Different from PLOT, we employ unbalanced OT to enhance the cooperation between global and local prompts by relaxing one of the equality constraints, which allows prompts to concentrate exclusively on the most relevant image patches.
@@ 5 Experiments
In this section, we conducted comprehensive experiments to numerically evaluate our FedOTP in the scenarios of heterogeneous data distribution.
@@ 5.1 Experimental Setup
Datasets and Data Heterogeneity. We evaluated the performance of FedOTP on nine public benchmark datasets with different types of heterogeneity, including label shift and feature shift. To investigate label shift, we selected two types of datasets. Following previous research [27, 26], we utilized five visual classification datasets to simulate datasets with limited samples: Food101 [3], DTD [11], Caltech101 [20], Flowers102 [54], and OxfordPets [56]. Referring to these datasets as the CLIP dataset for convenience, we utilized a Pathological setting by randomly allocating a distinct number of non-overlapping classes to each client. We also employed two image benchmark datasets: CIFAR-10, and CIFAR-100 [36]. We considered the Dirichlet distribution as introduced in [62, 39, 4] where datasets are partitioned randomly among clients using a symmetric Dirichlet distribution with $\alpha=0.3$. For feature shift, we utilized two datasets with multiple domains: DomainNet [58] with 6 domains, and Office-Caltech10 [25] with 4 domains. In line with prior studies [44, 67], each client participating in the federated learning system is assigned data from one of these distinct domains. The details of these dataset setup can be found in the Appendix Section B.1.
Baselines. We compared FedOTP with three kinds of baselines: (1) Local training methods: (i) Zero-shot CLIP [60] with hand-crafted text prompt templates; (ii) CoOp [78] with learnable prompt vectors trained on each client locally. (2) Existing prompt-based federated learning methods: (i) PromptFL [27] learning a unified prompt across clients; (ii) pFedPrompt [26] learning a shared prompt with personalized visual attention modules. (3) Four adapted methods derived from traditional PFL techniques, including PromptFL + FT [24], PromptFL+FedProx [42], PromptFL+FedPer [1] and PromptFL+FedAMP [32], as introduced in [26].
Implementation Details. To simulate federated learning in various scenarios, we consider the following two settings: (1) $n=10$ clients with a full $100\%$ partition, (2) $n=100$ clients with a $10\%$ partition. We employ SGD optimizer with a learning rate $lr=0.001$ and local epoch $R=5$ for CLIP datasets while $R=1$ for other cases. The communication round is set to $T=10$ for CLIP datasets with $10$ clients and $T=150$ for CIFAR-10/CIFAR-100 datasets with $100$ clients. We present the results using two representative backbones, ResNet50 [29] and ViT_B16 [15], defaulting to ViT_B16 if not explicitly specified. More implementation details can be found in the Appendix Section B.2.
@@ 5.2 Performance Evaluation
Model Evaluation on Label Shifts. We first measured the performance of FedOTP against baselines on datasets with label shifts. The experimental results on CLIP datasets and CIFAR-10/CIFAR-100 datasets are summarized in Table 1 and Table 3. For easy comparison, Table 1 reports results utilizing ResNet50 as the backbone, maintaining consistency with [26]. As shown in Table 1, our FedOTP outperforms state-of-the-art algorithms by a large margin across all datasets, which confirms the effectiveness of our Global-Local prompt cooperation mechanism to handle label shift scenarios. Remarkably, while both PromptFL+FedPer (which splits the learnable prompt vector into “base+personalized” vectors) and pFedPrompt (utilizing a shared prompt with a personalized attention module in the vision modal) experience significant declines when datasets are altered, FedOTP exhibits slight fluctuations. This verifies the robustness of our method across diverse scenarios. Table 3 shows the results of our FedOTP and benchmark methods on CIFAR-10/CIFAR-100 datasets under Dirichlet setting over $100$ clients with $10\%$ partition. Even in this scenario with Dirichlet settings and a large number of clients, FedOTP consistently outperforms the baseline methods, further highlighting the superiority of our approach.
@@ 10                   Calculate the cost matrix $C_{k}=1-G_{m}^{\top}H_{k}$ of each class;
CIFAR-10 [36] Image Classification 10 50,000 10,000 1
CIFAR-100 [36] Image Classification 100 50,000 10,000 1
@@ B.2 Implementation Details
All input images across datasets are resized to $224\times 224$ pixels and further divided into $14\times 14$ patches with a dimension of $768$. Regarding the hyperparameters for solving OT, we set the entropic regularization weight in Problem Eq. (8) as $\lambda=0.1$ for all datasets. The maximum iteration number $n$ for Algorithm 1 is set to $100$, and we implement early stopping when the absolute update value $\Delta_{v}$ is less than $0.001$. For the setting of learnable prompts, the length of prompt vectors $s$ is set to $16$ with a dimension of $512$, “end” token position, and “random” initialization. Batch sizes are set to $32$ for training and $100$ for testing. All experiments are conducted with Pytorch [57] on NVIDIA A40 GPUs.